OpenStudy (anonymous):

Hello I was wondering on how to evaluate the limit of: sqrt(6+h)-sqrt(6) -------------- h Lim h->0

OpenStudy (phi):

multiply top and bottom by sqrt(6+h)+ sqrt(6) and use the idea up top that you have (a-b)(a+b) = a^2 - b^2

OpenStudy (anonymous):

Ok let me try that

OpenStudy (anonymous):

Ok so now I got \[\frac{ (\sqrt{6+h}-\sqrt{6})^{2} }{ h(\sqrt{6+h}+\sqrt{6} }\] So then the square roots cancel where I get \[\frac { h}{ h(\sqrt{6+h}+\sqrt{6}) }\] So I cancel the "h" on the top and bottom to get: \[\frac{ 1 }{ \sqrt{6+h}+\sqrt{6} }\]

OpenStudy (phi):

the top should be \[ \left( \sqrt{6+h} \right)^2 - \left( \sqrt{6} \right)^2 \]

OpenStudy (anonymous):

@Matthew071 it is sqrt(6+h)+ sqrt(6), with a plus between them for the upper term as well

OpenStudy (anonymous):

Agreed and there fore the square roots cancel so on the top all i have remaining is H on the top, am I correct?

OpenStudy (phi):

now you have \[ \frac{1}{ \sqrt{6+h}+ \sqrt{6} } \] take the limit as h->0

OpenStudy (anonymous):

Ok so all I do is plug in h

OpenStudy (anonymous):

So I get: \[\frac{ 1 }{ 2\sqrt{6} }\]

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thank you so much! Have a wonderful day!